### v2

# DreamDrone

Abstract
We introduce DreamDrone, an innovative method for generating unbounded flythrough scenes from textual prompts. Central to our method is a novel feature-correspondence-guidance diffusion process, which utilizes the strong correspondence of intermediate features in the diffusion model. Leveraging this guidance strategy, we further propose an advanced technique for editing the intermediate latent code, enabling the generation of subsequent novel views with geometric consistency. Extensive experiments reveal that DreamDrone significantly surpasses existing methods, delivering highly authentic scene generation with exceptional visual quality. This approach marks a significant step in zero-shot perpetual view generation from textual prompts, enabling the creation of diverse scenes, including natural landscapes like oases and caves, as well as complex urban settings such as Lego-style street views. Our code is publicly available.
Abstract
In this supplementary material, we provide more comprehensive ablation studies, comparisons of visual results with text-to-video methods, and additional visual results of our method. Finally, we discuss the limitations of this approach.
Figure 1: Visualization results of DreamDrone. We visualize 30 continuous views for each prompt. The generated scenes are geometry-consistent across adjacent camera views and the details are added gradually when the camera is moving forward.
[…]
1 Introduction
Recent advances in vision and graphics have enabled the synthesis of multi-view consistent 3D scenes along extended camera trajectories, a capability previously explored in works like [17, 19, 3, 7]. This emerging task, termed perpetual view generation [19], involves synthesizing views from a flying camera along an arbitrarily long trajectory, starting from a single RGB image. Our study extends this concept, asking: Can we fly into scenes from any context, real or imagined, using textual prompts? This paper advances the novel concept of zero-shot, training-free perpetual view generation. This entails creating infinite scenes from textual prompts without optimization or fine-tuning on specific datasets. The challenge is considerable: as the camera moves forward, unseen regions must be seamlessly integrated, new details resolved as they approach the camera, all while maintaining photorealism and diversity. Existing research in video synthesis, image inpainting, and view synthesis does not fully address these requirements due to various limitations.
Recent text-to-video generation methods leverage advancements in diffusion models [2, 9, 40, 47, 14, 23] to extend their application to the temporal domain. These approaches show promise, but often require extensive training on large datasets, a costly and resource-intensive endeavor. T2V-0 [14] pioneered zero-shot text-to-video generation, producing notable results. However, limitations are evident as it can generate only a limited number of novel frames, and the quality degrades for longer videos. Additionally, these methods often overlook crucial video elements, such as scene geometry and camera movement. In contrast, text-to-3D methods like [18, 25, 30, 41, 44] utilize geometry to synthesize high-quality novel views but are constrained to limited camera motion and single-object view generation.
SceneScape [7], while conceptually similar to our work, diverges significantly in execution and capability. Its primary strategy involves outfilling regions during backward camera movement using 3D meshes. However, SceneScape encounters challenges with forward-moving cameras and in generating outdoor scenes. These issues stem from 1) difficulty in accurately filling irregular holes and preserving correspondence with original regions, 2) distortions in wrapped images remaining uncorrected, and 3) an inability to adequately detail areas enlarged by forward camera motion. Additionally, its dependence on mesh representations confines SceneScape to indoor scenarios, limiting its versatility. Our comparative analysis in Fig. 4 illustrates these limitations in SceneScape.
In this paper, we introduce DreamDrone, a novel method designed to generate infinite scenes from textual prompts. DreamDrone leverages off-the-shelf text-to-image diffusion models [34] and depth estimation models [32]. Central to our proposed method is a feature-correspondence-guidance diffusion process, specifically tailored to create geometry-consistent novel views during the denoising phase. Furthermore, we develop an innovative approach for editing the intermediate latent code, facilitating the generation of subsequent novel views. Complementing this, a cross-view self-attention module is employed to ensure consistent correspondence across adjacent views.
Our experiments demonstrate that the proposed DreamDrone effectively leads to high-quality and geometry-consistent scene generation, as evidenced in our results (see Fig. 1 and Figs. 3 and 4 in Sec. 4). Notably, despite other works relying on large-scale data for training, our method exhibits comparable, even superior performance. A significant advantage of DreamDrone is its versatility: it is adept not only at generating real-world scenarios but also shows promising capabilities in creating imaginative scenes. In summary, our key contributions are as follows:
  • We propose a novel approach for zero-shot text-to-perpetual view generation, capable of producing diverse indoor and outdoor scenes in various imaginative styles. This method achieves significant results without requiring fine-tuning or optimization on specific datasets.
  • We develop a novel feature-correspondence-guidance diffusion process to maintain geometric consistency in novel view generation. This approach involves a strategic modification of the diffusion model’s intermediate features, enhancing the fidelity and detail in each newly generated view.
  • Extensive experiments demonstrate that DreamDrone can generate infinite sequences, exemplified by 30 frames in Fig. 1. It handles a broad spectrum of scenarios, from realistic to fantastical, encompassing oases, caves, and urban scenes in styles like Lego or Van Gogh.
2 Related Works
Perpetual view generation.
InfNat [19], InfNat-0 [17], and DiffDreamer [3] use iterative training for long-trajectory perpetual view extrapolation. InfNat [19] pioneered the perpetual view generation task with a database for infinite 2D landscapes. InfNat-0 [17] adapted this to 3D, introducing a render-refine-repeat phase for novel views. DiffDreamer [3] improved consistency with image-conditioned diffusion models. However, trained on specific datasets, these methods lack robustness in new scenarios, like urban environments, and their non-text-based conditioning restricts generative capabilities. In very recent concurrent work, SceneScape [7] generates limited scenes when the camera is moving backward, which focuses on 3D-consistent ‘zoom out’ video generation. When the camera is moving forward, there is not any specific mechanism to add details and alleviate distortion. Moreover, SceneScape can only generate limited indoor scenes because they applies a mesh as a unified representation.
Text-to-3D generation.
Several text-to-3D generation methods [1, 27, 4, 29, 49] apply text-3D pair databases to learning a mapping function. However, supervised strategies remain challenging due to the lack of large-scale aligned text-3D pairs. CLIP-based [31] 3D generation methods [51, 28, 15, 13, 12] apply pre-trained CLIP model to create 3D objects by formulating the generation as an optimization problem in the image domain. Recent text-to-3D methods like [18, 24, 25, 44, 30] blend text-to-image diffusion models [34] with neural radiance fields [26] for training-free 3D object generation. Other approaches [37, 20, 21, 33] focus on novel view synthesis from a single image, often limited to single objects or small camera motion ranges. Text2room [10] generates 3D indoor scenes from text, but is confined to room meshes. Our DreamDrone diverges by creating endless, diverse indoor and outdoor scenes, in styles ranging from Lego to Van Gogh.
Text-to-video generation.
Generating videos from textual descriptions [23, 11, 9, 2, 36, 50, 43] poses significant challenges, primarily due to the scarcity of high-quality, large-scale text-video datasets and the inherent complexity in modeling temporal consistency and coherence. CogVideo [11] addresses this by incorporating temporal attention modules into the pre-trained text-to-image model CogView2 [6]. The video diffusion model [9] employs a space-time factorized U-Net, utilizing combined image and video data for training. Video LDM [2] adopts a latent diffusion approach for generating high-resolution videos. However, these methods typically do not account for the underlying 3D scene geometry in scene-related video generation, nor do they offer explicit control over camera movement. Additionally, their reliance on extensive training with large datasets can be prohibitively costly. While T2V-0 [14] introduced the concept of zero-shot text-to-video generation, its capability is limited to generating a small number of novel frames (up to 8 frames), with diminished quality in longer video sequences.
3 Method
Figure 2: Overview of DreamDrone. Starting from a real or generated RGBD ($I$, $D$) image at the current view, we apply DDIM backward steps to obtain intermediate latent code $x_{t_{1}}$ at timestep $t_{1}$ using a pre-trained U-Net model. A low-pass wrapping strategy is applied to generate latent code for the next novel view. A few more DDPM forward steps from timestep $t_{1}$ to $t_{2}$ are followed for enlarging the degree of freedom w.r.t. the wrapped latent code. In the reverse progress, we apply pre-trained U-Net to recover the novel view from $x_{t_{2}}^{\prime}$. The cross-view self-attention module and feature-correspondence guidance are applied to maintain consistency between $x_{t_{2}}$ and $x_{t_{2}}^{\prime}$. The right side shows the wrapped images and our generated novel view $I^{\prime}$. Our method greatly alleviates blurring, in-consistency, and distortion. The overall pipeline is zero-shot and training-free.
3.1 Preliminaries
In this paper, we implement our method based on the recent state-of-the-art text-to-image diffusion model (i.e. Stable Diffusion [34]). Stable diffusion is a latent diffusion model (LDM), which contains an autoencoder $\mathcal{D}(\mathcal{E}(\cdot)))$ and a U-Net [35] denoiser. Diffusion models are founded on two complementary random processes. The DDPM forward process, in which Gaussian noise is progressively added to the latent code of a clean image: $\boldsymbol{x}_{0}$:
$\boldsymbol{x}_{t}=\sqrt{\alpha_{t}}\boldsymbol{x}_{0}+\sqrt{1-\alpha_{t}}\boldsymbol{z},$ (1)
where $\boldsymbol{z}\sim\mathcal{N}(0,\mathbf{I})$ and $\left\{\alpha_{t}\right\}$ are the noise schedule.
The backward process is aimed at gradually denoising $\boldsymbol{x_{T}}$, where at each step a cleaner image is obtained. This process is achieved by a U-Net $\boldsymbol{\epsilon}_{\theta}$ that predicts the added noise $\boldsymbol{z}$. Each step of the backward process consists of applying $\boldsymbol{\epsilon}_{\theta}$ to the current $\boldsymbol{x}_{t}$, and adding a Gaussian noise perturbation to obtain a cleaner $\boldsymbol{x}_{t-1}$.
Classifier-guided DDIM sampling [5] aims to restore images from noise conditioned on the class label. Given the diffusion model $\boldsymbol{\epsilon}_{\theta}$, the latent code $\boldsymbol{x}_{t}$ at timestep $t$, the classifier $p_{\theta}(y|x_{t})$, and the gradient scale $s$, the sampling process for obtaining $\boldsymbol{x}_{t-1}$ is formulated as:
$\hat{\epsilon}=\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t})-\sqrt{\bar{\alpha}_{t-1}}\triangledown_{\boldsymbol{x}_{t}}\log p_{\phi}(y|\boldsymbol{x}_{t}),$ (2)
and
$\boldsymbol{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\cdot\frac{\boldsymbol{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{\epsilon}}{\sqrt{\bar{\alpha}_{t}}}+\sqrt{1-\bar{\alpha}_{t-1}}\hat{\epsilon},$ (3)
where $\alpha$ is the denoise schedule.
In the self-attention block of the U-Net, features are projected into queries $\mathit{\mathbf{Q}}$, keys $\mathit{\mathbf{K}}$, and values $\mathit{\mathbf{V}}$. The output of the block $\boldsymbol{o}$ is obtained by:
$\boldsymbol{o}=\mathbf{A}\mathbf{V},\;\;\textup{where}\;\mathbf{A}=\textup{Softmax}(\mathit{\mathbf{Q}}\mathit{\mathbf{K}}^{\top})$ (4)
The self-attention operation allows for long-range interactions between features.
3.2 Overview
In generating infinite scenes, the process involves two key steps. Firstly, the RGB for the forthcoming frame is calculated based on the RGBD data and camera extrinsics of the current frame. Secondly, to address the presence of black holes in the next frame’s image, inpainting [34, 22, 48] and outpainting [16, 46] techniques are employed for correction. Despite this approach, several challenges arise: 1) The irregular nature of the black holes makes it difficult for existing inpainting methods to fill these areas while maintaining semantic coherence with the rest of the image. 2) Inaccuracies in the depth data can cause distortions in certain image areas when wrapping. These distortions tend to accumulate over time with the movement of the camera, leading to a noticeable decline in image quality. 3) Additionally, as the image advances, it requires the enhancement of details in certain areas, a task at which current inpainting and outpainting methods are inadequate, particularly due to their lack of super-resolution capabilities.
To avoid generating distorted areas, ensure semantic consistency between adjacent frames, and enhance image detail during camera movement, we leverage the robustness and rich semantic capabilities of the denoise phase and decoder in the stable diffusion model. We edit the features of the diffusion based on depth and camera extrinsics, followed by denoising to produce a high-definition image for the subsequent frame. Our comprehensive design is detailed in Fig. 2. Initially, we obtain the latent code of the previous frame’s RGB image at timestep $t_{1}$ through the inverse diffusion process. We then wrap the current frame’s latent code using depth information and camera extrinsics. To maintain content and high-frequency detail consistency while wrapping the latent code, we introduce a low-pass wrapping strategy. Then several DDPM (Denoising Diffusion Probabilistic Models) processes are applied to increase the generative process’s flexibility, adding noise from timestep $t_{1}$ to $t_{2}$. Finally, we obtain the RGB image of the next frame through the denoising process. For ensuring semantic consistency between adjacent frames, we utilize a cross-frame self-attention strategy, injecting the key and value from the previous frame’s UNet network into the next frame. Furthermore, we propose a feature-correspondence-guidance strategy to influence and maintain semantic consistency between adjacent frames during the denoise latent code process. Our entire pipeline requires only a pre-trained text-to-image diffusion model and a depth estimation model, eliminating the need for any additional training or fine-tuning.
3.3 Wrapping latent codes
Algorithm 1 Low-pass latent code wrapping
$\boldsymbol{x}_{t}$$\triangleright$ latent code at timestep $t$ of current view $c$
1:$F(\boldsymbol{x}_{t})\leftarrow FFT(\boldsymbol{x}_{t})$triangleright Apply Fast Fourier Transform
2:Split $F(\boldsymbol{x}_{t})$ into $F_{low}$ and $F_{high}$ using threshold $\sigma$
3:$\boldsymbol{x}_{t}^{low}\leftarrow IFFT(F_{low})$triangleright Inverse FFT on low-frequency component
4:$\boldsymbol{x}_{t}^{low-wrapped}\leftarrow Wrap(\boldsymbol{x}_{t}^{low})$triangleright Wrap the low-frequency content
5:$F_{wrapped}\leftarrow FFT(\boldsymbol{x}_{t}^{low-wrapped})$triangleright FFT on wrapped content
6:$F^{\prime}\leftarrow F_{wrapped}+F_{high}$triangleright Combine low-frequency of wrapped content with high-frequency of original code
[…]
The results in the right side of Fig. 2 reveal that directly wrapping images based on camera intrinsics $K$, extrinsics ($R,t$), and depth information leads to regions of distortion in the images. Additionally, the use of inpainting [34, 22, 48] and outpainting [16, 46] models to fill these gaps does not achieve satisfactory outcomes. In pursuit of generating photo-realistic images, and considering the robustness of the denoising process and decoder in the stable diffusion model, we opt to edit the latent code corresponding to timestep $t$. PnP [39] and DIFT [38] have shown that the features of diffusion possess strong semantic information, with semantic parts being shared across images at each step. The simplest method for wrapping the latent code follows the same approach as wrapping the image. The only difference between wrapping the latent code and wrapping the image is a slight modification in the camera intrinsics; this entails scaling the camera intrinsics proportionally based on the different resolutions of the image and latent code.
The overall procedure for wrapping the latent code is illustrated in Alg. 1. Initially, a latent code ${x}_{t}$ is obtained and transformed via Fast Fourier Transform (FFT) to $F(\boldsymbol{x}_{t})$. This is divided into low-frequency $F_{\text{low}}$ and high-frequency $F_{\text{high}}$ components, segregated at threshold $\lambda$. The key step involves wrapping the Inverse FFT (IFFT) processed low-frequency component $x_{t}^{\text{low}}=\text{IFFT}(F_{\text{low}})$, forming $\boldsymbol{x}_{t}^{low-wrapped}$. Merging $\text{FFT}(\boldsymbol{x}_{t}^{low-wrapped})$ with $F_{\text{high}}$, we obtain $F_{\text{new}}$, from which the final latent code $\boldsymbol{x}_{t}^{\prime}=\text{IFFT}(F^{\prime})$ is reconstructed. This approach efficiently preserves high-frequency details, enabling high-fidelity scene generation aligned with text prompts.
3.4 Feature-correspondence-guidance design
After obtaining the latent code $x_{t}^{\prime}$ corresponding to the next frame, we employ the DDPM (Denoising Diffusion Probabilistic Models) method to increase the degrees of freedom of the latent code, enabling the generation of richer image details. However, increasing freedom introduces a challenge: the correlation between frames. An unconstrained diffusion denoising process can result in poor semantic correlation between adjacent frames. To address this, we propose a feature-correspondence guidance strategy with a cross-view self-attention mechanism. We introduce these approaches in detail below.
Cross-view self-attention.
To maintain consistency between the generated result and the original image, inspired by recent image and video editing works [39, 45, 8, 42], we modify the process of the self-attention module of U-Net when denoising the latent code $\boldsymbol{x}_{t}^{\prime}$. Specifically, we denoise the views for the current and next view together. The key and value of the self-attention modules from the next view are replaced by that of the current view. To be specific, for obtaining the original view, the self-attention module is defined the same as Eq. 4. The modified cross-view self-attention for generating a novel view is defined as:
$\boldsymbol{o}^{\prime}=\mathbf{A^{\prime}}\mathbf{V},\;\;\textup{where}\;\mathbf{A}^{\prime}=\textup{Softmax}(\mathbf{Q}^{\prime}\mathbf{K}^{\top}),$ (5)
where $\mathbf{Q}^{\prime}$, $\mathbf{A}^{\prime}$, and $\boldsymbol{o}^{\prime}$ are query, attention matrix, and output features for the novel views. $\mathbf{K}$ and $\mathbf{V}$ are injected keys and values obtained from the self-attention module for generating the original view.
Feature-correspondence guidance.
Maintaining geometry consistency between adjacent views using the cross-view self-attention mechanism presents challenges, especially in preserving high-frequency details as the camera moves forward. The recent DIFT study [38] highlights the potential of using intermediate features of diffusion models for accurate point-to-point image matching. Additionally, the concept of vanilla classifier guidance [5] steers the diffusion sampling process using pre-trained classifier gradients towards specific class labels. Building on these ideas, we integrate feature correspondence guidance into the DDIM process to enhance consistency between adjacent views, addressing the challenge of detail preservation in dynamic scenes.
Specifically, we obtain the features of the current and next view at each timestep $t$ of the DDIM process and calculate the cosine distance between the wrapped original features and features from the next novel views:
$\mathcal{L}_{sim}^{t}=\frac{1-\cos\left[\mathrm{Wrap}(f_{t}),f_{t}^{\prime}\right]}{2},$ (6)
where $f_{t}$ and $f_{t}^{\prime}$ are diffusion features extracted from pre-trained U-Net $\boldsymbol{\epsilon}_{\theta}$ at timestep $t$ and $\mathrm{Wrap}$ is the wrapping functions. The lower $\mathcal{L}_{sim}^{t}$, the higher the similarity.
We further introduce the similarity score $\mathcal{L}_{sim}$ to the DDIM sampling process, for generating novel views with geometry consistency. The predicted noise $\hat{\epsilon}$ is formulated as:
$\hat{\epsilon}=\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t})-\lambda\sqrt{\bar{\alpha}_{t-1}}\triangledown_{\boldsymbol{x}_{t}}\mathcal{L}_{sim}^{t},$ (7)
where $\lambda$ is the constant hyper-parameter and latent code $\boldsymbol{x}_{t-1}$ is calculated by Eq. 3
4 Experiments
Figure 3: Qualitative comparisons of InfNat-0 [17] and ours. We provide four scene images with various styles and categories as start points and ask models to fly through the images. 10 frames are illustrated for each starting point.
Implementation details.
We take Stable Diffusion [34] with the pre-trained weights from version 2.1https://huggingface.co/stabilityai/stable-diffusion-2-1-base as the basic text-to-image diffusion and MiDas [32] with weights $\mathrm{dpt\_beit\_large\_512}$https://github.com/isl-org/MiDaS. The overall diffusion timesteps is $1000$. We wrap the latent code at timestep $t_{1}$=$21$ and add more degrees of noise to timestep $t_{2}$=$441$. The threshold $\sigma$ for low-pass wrapping is $20$ and the hyper-parameter $\lambda$ for feature-correspondence guidance is $300$. Due to the page limit, please refer to the supplementary material (supp.) for details.
Baselines.
We compare against 1) 2 supervised methods for perpetual view generation: InfNat [19] and InfNat-0 [17]. 2) one text-conditioned scene generation: SceneScape [7]. 3) 2 supervised methods for text-to-video generation: CogVideo [11] and VideoFusion [23]. 4) one method for zero-shot text-to-video generation: T2V-0 [14].
Evaluation Metrics.
The perpetual view generation is evaluated both quantitatively and qualitatively. We evaluate CLIP [31] score, which indicates text-scene alignment for quantitative comparisons. CogVideo [11], VideoFusion [23], SceneScape [7], and T2V-0 [14] are all engaged in text-conditioned generation tasks. We generated 50 scene-related text prompts using GPT-4https://openai.com/gpt-4 and then created videos using each of the three methods. For the InfNat [19] and InfNat-0 [17] methods, we used Stable Diffusion to generate the initial frame, followed by subsequent frame generation based on this initial frame. We calculated the distance between each generated frame and the text embedding, known as the CLIP score. A high average CLIP score indicates not only that the generated images are more aligned with the corresponding prompts but also that they consistently maintain high quality. Considering that the InfNat [19] and InfNat-0 [17] methods might gradually generate other scenes, we provided 10 very general prompts such as ‘an image of the landscape’ and ‘an image of the mountain’ for these methods, and then selected the highest CLIP score as the CLIP score for the current frame.
4.1 Quantitative comparison
Figure 4: Qualitative comparisons of SceneScape [7], T2V-0 [14], and our DreamDrone. We visualize 20 continuous frames for each textual prompt. As the camera flies, our method generates geometry-consistent scene sequences.
The quantitative comparison results are shown in Tab. 1. It is evident that the CLIP scores for the InfNat and InfNat-0 methods are low, indicating poor robustness. These methods are not sufficiently versatile to handle infinite scene generation for any scenario, as also illustrated by the visual results in Fig. 3. CogVideo and VideoFusion, which are supervised text-to-video generation methods, struggle to generate first-person perspective landscapes due to limitations of text-video training datasets. Additionally, their CLIP scores decrease with an increasing number of frames, such as 32 frames, which is corroborated by the visual results in supp. T2V-0 [14], SceneScape [7], and our method, both based on the pre-trained stable diffusion model and training-free, have relatively higher CLIP scores. However, as the number of frames increases, SceneScape and T2V-0’s CLIP scores decrease, whereas ours remains constant (please refer to Fig. 4 for qualitative comparisons). We attribute this to our low-pass wrapping strategy, which provides better content guidance for the generated images, and the introduction of feature-correspondence guidance, ensuring the coherence of adjacent frames and thus maintaining a higher CLIP score.
Table 1: Quantitative comparison results. We calculate the average CLIP score [31] across various video lengths.
8 frames 16 frames 32 frames
InfNat [19] 0.125 0.123 0.118
InfNat-0 [17] 0.128 0.125 0.122
SceneScape [7] 0.263 0.238 0.174
CogVideo [11] 0.255 0.249 0.241
VideoFusion [23] 0.281 0.283 0.272
T2V-0 [14] 0.312 0.305 0.287
Ours 0.320 0.318 0.319
4.2 Qualitative comparison
In our comparison with InfNat-0 [17] (Fig. 3), focusing on various scenes including coastlines, rivers, Van Gogh-style landscapes, and city streetscapes, we identified four main differences: Firstly, InfNat-0 shows proficiency in coastline scenes, a reflection of its training data, but our training-free DreamDrone surpasses it in later frames due to InfNat-0’s cumulative errors over time. Secondly, in natural scenes with closer objects, InfNat-0’s flawed generation becomes more apparent, whereas our method maintains consistency. Thirdly, InfNat-0’s limited approach to gap filling leads to poor performance in stylized scenes, in contrast to DreamDrone which preserves high-frequency details and frame correspondence. Finally, in urban environments, InfNat-0 struggles significantly, while DreamDrone achieves realistic and geometry-consistent views, demonstrating its versatility across varied scenarios.
T2V-0 [14] introduces unsupervised text-conditioned video generation using stable diffusion. SceneScape [7] focuses on ‘zoom out’ effects during backward camera movement. However, as seen in Fig. 4, both methods have limitations. SceneScape struggles with outdoor scenes and forward camera movement, leading to blurred and distorted results after 8 steps due to its reliance on a pre-trained inpainting model. T2V-0 displays a decline in quality beyond the third frame in complex environments like Lego-style cities, likely from its latent code editing approach that compromises frame continuity and geometric consistency. Conversely, our DreamDrone excels across various scenes. It maintains detail, continuity, and quality in advancing camera scenarios, evident in even simpler landscapes like mountains where T2V-0 and SceneScape cannot effectively portray dynamic elements like cloud movement. Our approach ensures the preservation of fine details such as shadows and sunlight, creating a more dynamic and realistic video experience. Please refer to supp. for more quantitative comparisons.
Our task bears similarities to text-to-video generation, we further provide qualitative comparisons with VideoFusion [23], one of the state-of-the-art methods for text-to-video generation. Due to the page limit, please refer to supp. for detailed comparisons.
4.3 Ablation studies
Figure 5: Ablation study on the key components of our method. We perform ablation studies by disabling the key components of our method: feature-correspondence guidance, low-pass wrapping strategy, and cross-view self-attention module.
We perform ablation studies on 1) low-pass feature wrapping, 2) cross-view self-attention module, and 3) feature-correspondence guidance in Fig. 5. Without feature-correspondence guidance and cross-view attention, even if we perform the corresponding wrapping at the feature level, it is challenging to ensure strong consistency between generated adjacent frames ($1^{st}$ row). Direct wrapping of the latent code leads to increasingly blurred images and a loss of high-frequency details. However, the use of our low-pass wrapping strategy effectively ensures that the generated images retain high-definition details ($2^{nd}$ row). Without the guidance of the feature-correspondence guidance mechanism or the presence of a cross-view self-attention module, it is difficult to guarantee the geometric consistency of adjacent frames ($3^{rd}$ and $4^{th}$ rows). The components in our proposed pipeline enhance each other, ensuring that each generated frame is not only rich in detail but also maintains good consistency with adjacent frames. Please refer to supp. for more detailed ablation studies.
5 Conclusion
In this work, we pioneer the concept of zero-shot, training-free text-to-perpetual view generation, introducing DreamDrone for generating geometry-consistent scenes. We harness the robust capabilities of a pre-trained text-to-image diffusion model and develop a novel feature-correspondence-guidance diffusion process for creating novel views. Our quantitative and qualitative analyses underscore DreamDrone’s superiority over other state-of-the-art methods. We anticipate that our approach will inspire further innovation in AI-Generated Content (AIGC) and make advanced research accessible to a broader spectrum of researchers.
Supplementary Material
6 Implementation details
We take Stable Diffusion [34] with the pre-trained weights from version 2.1https://huggingface.co/stabilityai/stable-diffusion-2-1-base as the basic text-to-image diffusion and MiDas [32] with weights $\mathrm{dpt\_beit\_large\_512}$https://github.com/isl-org/MiDaS. The overall diffusion timesteps is $1000$. We wrap the latent code at timestep $t_{1}$=$21$ and add more degrees of noise to timestep $t_{2}$=$441$. The threshold $\sigma$ for low-pass wrapping is $20$ and the hyper-parameter $\lambda$ for feature-correspondence guidance is $300$. We conducted the experiments on Titan-RTX GPU. The generated speed is roughly 30 seconds per frame.
7 Ablation studies
In this section, we conduct our ablation studies on each proposed module. First, we explore the impact of wrapping the latent code at timestep $t_{1}$ in our pipeline on the generation results, with comparative results shown in Fig. 6. According to the principles of diffusion, a smaller timestep means that the corresponding latent code has less noise, thus retaining more complete content and detailed information. When we wrap the latent code at a smaller timestep (for example,$t_{1}=1$ or $21$), it is observed that, with the camera’s movement, DreamDrone still maintains its complete content information quite well. As the timestep increases, with more noise, the content information gets somewhat disrupted, and wrapping the latent code then does not yield as favorable results. This experiment demonstrates that the latent code corresponding to a smaller timestep possesses richer geometry and content information.
Figure 6: Ablation study on the wrapping timestep $t_{1}$.
Next, we investigate how editing the latent code in different ways can enhance the generation results. We set various thresholds for the low-pass wrapping module, where a smaller threshold indicates the retention of less low-frequency information and more high-frequency information. The comparative results are detailed in Fig. 7. The first and second rows display the generation results corresponding to smaller thresholds. It can be observed that setting an excessively small threshold retains too much ineffective or noisy high-frequency information, leading to generated images with an unrealistic abundance of high-frequency details. Conversely, setting the threshold too high, or removing the low-pass wrapping module, results in a lack of retained detail and very blurry images.
Figure 7: Ablation study on the threshold of the low-pass wrapping module $\sigma$.
After obtaining the wrapped latent code, we next investigate the impact of setting $t_{2}$ on the generation results. The size of $t_{2}$ influences the freedom in the denoising process; a larger $t_{2}$ means adding more noise to the wrapped latent code, thereby gaining higher degrees of freedom. The generation results corresponding to different $t_{2}$ settings are shown in Fig. 8. A smaller $t_{2}$ (for example, $t_{2}=161$ or 241) results in insufficient noise added to the wrapped latent code, leading to increasingly blurred generated images. Setting a larger $t_{2}$ produces clearer images with richer details, but an excessively high degree of freedom can decrease the geometric consistency between adjacent frames, adversely affecting their coherence.
Figure 8: Ablation study on the denoising timestep $t_{2}$.
To ensure geometric consistency between adjacent frames during the denoising process, we introduced a feature-correspondence guidance strategy. Different guidance weights lead to significant variations in generation results, as detailed in Fig. 9. When feature-correspondence guidance is not introduced (i.e., $\lambda=0$), the generated image sequence maintains content consistency, but some details may not be consistent. The introduction of feature-correspondence guidance ensures the consistency of details between adjacent frames; however, excessively high guidance weights can impact the realism of the generated images.
Figure 9: Ablation study on the weight parameter of feature-correspondence guidance $\lambda$.
8 Additional qualitative comparisons
Figure 10: Qualitative comparisons of VideoFusion [23] and our DreamDrone.
Our task bears similarities to text-to-video generation, with the key difference being that text-to-video generation cannot be controlled by camera pose, and the quality significantly diminishes as the number of generated frames increases. VideoFusion [23], one of the state-of-the-art methods for video generation tasks, has been visually compared with our method, which is illustrated in Fig. 10. It is evident that VideoFusion’s generated results become blurry with an increase in frame count, and the effect of camera movement is less pronounced. In contrast, our method not only generates high-quality continuous scenes but also ensures geometric consistency between frames, clearly conveying the camera’s forward movement. Generating scenes in constrained environments like caves is more challenging. VideoFusion does not perform well under such prompts, whereas our method effectively demonstrates the effect of the camera advancing forward.
9 More visualization results
In this section, we provide more visualization results. Please refer to Figs. 11 and 12 for details. Moreover, please watch our attached videos for a comprehensive evaluation of our proposed DreamDrone.
Figure 11: Visualization results of our DreamDrone.
Figure 12: Visualization results of our DreamDrone.
10 Limitation
Given a prompt, our method can infinitely extend a scene without any training or fine-tuning. However, there are some limitations to our approach. Firstly, as our method is zero-shot and training-free, even with the introduction of feature-correspondence guidance and cross-frame self-attention modules, the correspondence of high-frequency details between adjacent frames is not yet perfect. Secondly, our method heavily relies on the accuracy of depth estimation. Although the stable diffusion model exhibits some robustness, for scenes with special styles, the entirely incorrect depth information leads to unsatisfactory generation results. We plan to address these shortcomings in our future work.

### v3

@@
-DreamDrone
+DreamDrone: Text-to-Image Diffusion Models are Zero-shot Perpetual View Generators
 Abstract
-We introduce DreamDrone, an innovative method for generating unbounded flythrough scenes from textual prompts.
-Central to our method is a novel feature-correspondence-guidance diffusion process, which utilizes the strong correspondence of intermediate features in the diffusion model.
-Leveraging this guidance strategy, we further propose an advanced technique for editing the intermediate latent code, enabling the generation of subsequent novel views with geometric consistency.
-Extensive experiments reveal that DreamDrone significantly surpasses existing methods, delivering highly authentic scene generation with exceptional visual quality.
-This approach marks a significant step in zero-shot perpetual view generation from textual prompts, enabling the creation of diverse scenes, including natural landscapes like oases and caves, as well as complex urban settings such as Lego-style street views.
-Our code is publicly available.
-Abstract
-In this supplementary material, we provide more comprehensive ablation studies, comparisons of visual results with text-to-video methods, and additional visual results of our method.
-Finally, we discuss the limitations of this approach.
+We introduce DreamDrone, a novel zero-shot and training-free pipeline for generating unbounded flythrough scenes from textual prompts.
+Different from other methods that focus on warping images frame by frame, we advocate explicitly warping the intermediate latent code of the pre-trained text-to-image diffusion model for high-quality image generation and generalization ability.
+To further enhance the fidelity of the generated images, we also propose a feature-correspondence-guidance diffusion process and a high-pass filtering strategy to promote geometric consistency and high-frequency detail consistency, respectively.
+Extensive experiments reveal that DreamDrone significantly surpasses existing methods, delivering highly authentic scene generation with exceptional visual quality, without training or fine-tuning on datasets or reconstructing 3D point clouds in advance.
 Figure 1: Visualization results of DreamDrone.
-We visualize 30 continuous views for each prompt.
-The generated scenes are geometry-consistent across adjacent camera views and the details are added gradually when the camera is moving forward.
-[…]
+Given a single scene image and the textual description, our approach generates novel views corresponding to user-defined camera trajectory, without fine-tuning on any dataset or reconstructing the 3D point cloud in advance.
 1 Introduction
-Recent advances in vision and graphics have enabled the synthesis of multi-view consistent 3D scenes along extended camera trajectories, a capability previously explored in works like [17, 19, 3, 7].
-This emerging task, termed perpetual view generation [19], involves synthesizing views from a flying camera along an arbitrarily long trajectory, starting from a single RGB image.
-Our study extends this concept, asking: Can we fly into scenes from any context, real or imagined, using textual prompts?
-This paper advances the novel concept of zero-shot, training-free perpetual view generation.
-This entails creating infinite scenes from textual prompts without optimization or fine-tuning on specific datasets.
-The challenge is considerable: as the camera moves forward, unseen regions must be seamlessly integrated, new details resolved as they approach the camera, all while maintaining photorealism and diversity.
-Existing research in video synthesis, image inpainting, and view synthesis does not fully address these requirements due to various limitations.
-Recent text-to-video generation methods leverage advancements in diffusion models [2, 9, 40, 47, 14, 23] to extend their application to the temporal domain.
-These approaches show promise, but often require extensive training on large datasets, a costly and resource-intensive endeavor.
-T2V-0 [14] pioneered zero-shot text-to-video generation, producing notable results.
-However, limitations are evident as it can generate only a limited number of novel frames, and the quality degrades for longer videos.
-Additionally, these methods often overlook crucial video elements, such as scene geometry and camera movement.
-In contrast, text-to-3D methods like [18, 25, 30, 41, 44] utilize geometry to synthesize high-quality novel views but are constrained to limited camera motion and single-object view generation.
-SceneScape [7], while conceptually similar to our work, diverges significantly in execution and capability.
-Its primary strategy involves outfilling regions during backward camera movement using 3D meshes.
-However, SceneScape encounters challenges with forward-moving cameras and in generating outdoor scenes.
-These issues stem from 1) difficulty in accurately filling irregular holes and preserving correspondence with original regions, 2) distortions in wrapped images remaining uncorrected, and 3) an inability to adequately detail areas enlarged by forward camera motion.
-Additionally, its dependence on mesh representations confines SceneScape to indoor scenarios, limiting its versatility.
-Our comparative analysis in Fig.
-4 illustrates these limitations in SceneScape.
-In this paper, we introduce DreamDrone, a novel method designed to generate infinite scenes from textual prompts.
-DreamDrone leverages off-the-shelf text-to-image diffusion models [34] and depth estimation models [32].
-Central to our proposed method is a feature-correspondence-guidance diffusion process, specifically tailored to create geometry-consistent novel views during the denoising phase.
-Furthermore, we develop an innovative approach for editing the intermediate latent code, facilitating the generation of subsequent novel views.
-Complementing this, a cross-view self-attention module is employed to ensure consistent correspondence across adjacent views.
-Our experiments demonstrate that the proposed DreamDrone effectively leads to high-quality and geometry-consistent scene generation, as evidenced in our results (see Fig.
-1 and Figs.
-3 and 4 in Sec.
+Recent advances in vision and graphics have enabled the synthesis of multi-view consistent 3D scenes along extended camera trajectories [18, 20, 3, 7].
+This emerging task, termed perpetual view generation [20], involves synthesizing views from a flying camera along an arbitrarily long trajectory, starting from a single RGBD image.
+Previous methodologies predominantly engage in warping images frame by frame with traditional 3D geometric knowledge when given RGBD images and subsequent camera extrinsic.
+However, this operation often leads to blurriness and distortion in images, which arises from inaccurate interpolation, the mismatch between discrete pixels and continuous transformations, and inaccurate depth data.
+Moreover, such blurriness and distortion tend to amplify with the accumulation of warp operations.
+To further alleviate the errors caused by frame-by-frame warp operations, two primary paths have been proposed.
+i) Some methods [18, 20, 3] try to train a refiner on natural scene datasets.
+The advantage of this frame-by-frame approach is that it allows for arbitrary changes in camera trajectory during the scene generation process, offering users a higher degree of freedom and enabling infinite generation.
+However, this training-based method can only be used in natural scenes and cannot be generalized to arbitrary indoor/outdoor scenes or scenes of various styles.
+ii) Another solution is to first reconstruct the 3D scene model using text prompts, then render 2D RGB images according to the camera trajectory [7, 52, 10].
+Although this solution yields more coherent 2D image sequences, the quality of the rendered images highly depends on the quality of the 3D scene model.
+This method cannot guarantee good rendering effects from every viewpoint.
+Additionally, since this method requires the reconstruction of 3D point clouds, it cannot achieve "infinite" scene generation in the same way as the frame-by-frame strategy.
+In this paper, we advocate that a more general and flexible perpetual view generation pipeline should possess the following capabilities:
+i) are versatile across diverse scenes, including indoor and outdoor scenes, as well as scenes depicted in various styles; ii) allow users to interactively control the camera trajectory during the process of scene generation, while ensuring the high quality of the generated images and the semantic consistency between adjacent frames; and iii) enable seamless transitions from one scene to another.
+To this end, we introduce DreamDrone, a novel zero-shot, training-free, infinite scene generation pipeline from text prompts, which does not require any optimization or fine-tuning on any dataset.
+A core principle of our approach is to warp the latent code of a pre-trained text-to-image diffusion model rather than the frames, enriching it with temporal and geometric consistency.
+To be specific, given RGBD image $I$ of the current view and camera rotation $\mathcal{R}$ and translation $\mathcal{T}$ for the next view (which is interactively defined by users), we first obtain the latent code $x_{t_{1}}$ of the diffusion model at timestep $t_{1}$, warp it to latent code of the next view $x^{\prime}_{t_{1}}$ based on $\mathcal{R}$ and $\mathcal{T}$, and denoise $x^{\prime}_{t_{1}}$ to the image $I^{\prime}$ of the next view.
+To ensure geometry consistency across adjacent views, we propose a novel feature-correspondence-guidance diffusion process when denoising from $x^{\prime}_{t_{1}}$ to image at the next view $I^{\prime}$.
+Moreover, we propose a novel high-pass filter mechanism when warping the latent code $x_{t_{1}}$, for preserving high-frequency details across adjacent views.
+Our experiments demonstrate that the proposed DreamDrone effectively leads to high-quality and geometry-consistent scene generation.
+Quantitative and qualitative results demonstrate our comparable, even superior performance compared with other training-based and training-free methods from the aspects of temporal consistency and image quality.
+Moreover, the significant advantage of DreamDrone is its versatility: it is adept not only at generating real-world scenarios but also shows promising capabilities in creating imaginative scenes.
+Additionally, users can interactively control the camera trajectory (Fig.
+5) and shuttle from one scene to another (Fig.
 4).
-Notably, despite other works relying on large-scale data for training, our method exhibits comparable, even superior performance.
-A significant advantage of DreamDrone is its versatility: it is adept not only at generating real-world scenarios but also shows promising capabilities in creating imaginative scenes.
-In summary, our key contributions are as follows:
-  • We propose a novel approach for zero-shot text-to-perpetual view generation, capable of producing diverse indoor and outdoor scenes in various imaginative styles.
-This method achieves significant results without requiring fine-tuning or optimization on specific datasets.
-  • We develop a novel feature-correspondence-guidance diffusion process to maintain geometric consistency in novel view generation.
-This approach involves a strategic modification of the diffusion model’s intermediate features, enhancing the fidelity and detail in each newly generated view.
-  • Extensive experiments demonstrate that DreamDrone can generate infinite sequences, exemplified by 30 frames in Fig.
-1.
-It handles a broad spectrum of scenarios, from realistic to fantastical, encompassing oases, caves, and urban scenes in styles like Lego or Van Gogh.
+Our contributions are summarized as follows:
+  • To our best knowledge, we are the first attempt to generate novel views by explicitly warping the latent code of the pre-trained diffusion model.
+  • A novel feature-correspondence-guidance diffusion process is proposed to enforce geometry consistency across adjacent views.
+Moreover, a high-pass filtering strategy is introduced to preserve high-frequency details for novel views.
+  • Extensive experiments demonstrate that our method generates high-quality and geometry-consistent novel views for any scene, from realistic to fantastical.
+More interestingly, our method realizes the scene shuttle, i.e., travels from one scene to another when the user controls the camera trajectory.
 2 Related Works
 Perpetual view generation.
+Perpetual view generation extrapolates unseen content outside a single image.
 InfNat [20], InfNat-0 [18], and DiffDreamer [3] use iterative training for long-trajectory perpetual view extrapolation.
@@
 DiffDreamer [3] improved consistency with image-conditioned diffusion models.
-However, trained on specific datasets, these methods lack robustness in new scenarios, like urban environments, and their non-text-based conditioning restricts generative capabilities.
-In very recent concurrent work, SceneScape [7] generates limited scenes when the camera is moving backward, which focuses on 3D-consistent ‘zoom out’ video generation.
-When the camera is moving forward, there is not any specific mechanism to add details and alleviate distortion.
-Moreover, SceneScape can only generate limited indoor scenes because they applies a mesh as a unified representation.
+However, these methods lack robustness in complex and urban environments.
+In very recent concurrent work, SceneScape [7] and WonderJourney [52] firstly generate 3D point cloud for scene by zoom-out and inpainting strategy.
+2D image sequences are further rendered based on the reconstructed 3D point cloud.
+However, the accuracy of the 3D model critically impacts performance, particularly with novel camera trajectories.
 Text-to-3D generation.
-Several text-to-3D generation methods [1, 27, 4, 29, 49] apply text-3D pair databases to learning a mapping function.
+Several text-to-3D generation methods [1, 28, 4, 30, 54, 15] apply text-3D pair databases to learning a mapping function.
 However, supervised strategies remain challenging due to the lack of large-scale aligned text-3D pairs.
 CLIP-based [33] 3D generation methods [56, 29, 16, 13, 12] apply pre-trained CLIP model to create 3D objects by formulating the generation as an optimization problem in the image domain.
-Recent text-to-3D methods like [18, 24, 25, 44, 30] blend text-to-image diffusion models [34] with neural radiance fields [26] for training-free 3D object generation.
+Recent text-to-3D methods like [19, 25, 26, 47, 31, 38, 50] blend text-to-image diffusion models [36] with neural radiance fields [27] for training-free 3D object generation.
 Other approaches [41, 21, 22, 35] focus on novel view synthesis from a single image, often limited to single objects or small camera motion ranges.
-Text2room [10] generates 3D indoor scenes from text, but is confined to room meshes.
-Our DreamDrone diverges by creating endless, diverse indoor and outdoor scenes, in styles ranging from Lego to Van Gogh.
+Text2room [10] generates 3D indoor scenes from text prompts, but is confined to room meshes.
 Text-to-video generation.
-Generating videos from textual descriptions [23, 11, 9, 2, 36, 50, 43] poses significant challenges, primarily due to the scarcity of high-quality, large-scale text-video datasets and the inherent complexity in modeling temporal consistency and coherence.
+Generating videos from textual descriptions [24, 11, 9, 2, 39, 55, 46, 40] poses significant challenges, primarily due to the scarcity of high-quality, large-scale text-video datasets and the inherent complexity in modeling temporal consistency and coherence.
 CogVideo [11] addresses this by incorporating temporal attention modules into the pre-trained text-to-image model CogView2 [6].
@@
 Additionally, their reliance on extensive training with large datasets can be prohibitively costly.
-While T2V-0 [14] introduced the concept of zero-shot text-to-video generation, its capability is limited to generating a small number of novel frames (up to 8 frames), with diminished quality in longer video sequences.
+While T2V-0 [14] introduced the concept of zero-shot text-to-video generation, its capability is limited to generating a small number of novel frames, with diminished quality in longer video sequences.
 3 Method
-Figure 2: Overview of DreamDrone.
-Starting from a real or generated RGBD ($I$, $D$) image at the current view, we apply DDIM backward steps to obtain intermediate latent code $x_{t_{1}}$ at timestep $t_{1}$ using a pre-trained U-Net model.
-A low-pass wrapping strategy is applied to generate latent code for the next novel view.
-A few more DDPM forward steps from timestep $t_{1}$ to $t_{2}$ are followed for enlarging the degree of freedom w.r.t.
-the wrapped latent code.
-In the reverse progress, we apply pre-trained U-Net to recover the novel view from $x_{t_{2}}^{\prime}$.
-The cross-view self-attention module and feature-correspondence guidance are applied to maintain consistency between $x_{t_{2}}$ and $x_{t_{2}}^{\prime}$.
-The right side shows the wrapped images and our generated novel view $I^{\prime}$.
-Our method greatly alleviates blurring, in-consistency, and distortion.
-The overall pipeline is zero-shot and training-free.
+We formulate the task of perpetual view generation as follows: given a starting image $I$, we generate the next view image $I^{\prime}$ corresponding to an arbitrary camera pose $\left\{\mathcal{R},\mathcal{T}\right\}$, where the camera pose can be specified or via user’s control.
 3.1 Preliminaries
-In this paper, we implement our method based on the recent state-of-the-art text-to-image diffusion model (i.e.
+We implement our method based on the recent state-of-the-art text-to-image diffusion model (i.e.
 Stable Diffusion [36]).
@@
 Diffusion models are founded on two complementary random processes.
-The DDPM forward process, in which Gaussian noise is progressively added to the latent code of a clean image: $\boldsymbol{x}_{0}$:
-$\boldsymbol{x}_{t}=\sqrt{\alpha_{t}}\boldsymbol{x}_{0}+\sqrt{1-\alpha_{t}}\boldsymbol{z},$ (1)
-where $\boldsymbol{z}\sim\mathcal{N}(0,\mathbf{I})$ and $\left\{\alpha_{t}\right\}$ are the noise schedule.
-The backward process is aimed at gradually denoising $\boldsymbol{x_{T}}$, where at each step a cleaner image is obtained.
-This process is achieved by a U-Net $\boldsymbol{\epsilon}_{\theta}$ that predicts the added noise $\boldsymbol{z}$.
-Each step of the backward process consists of applying $\boldsymbol{\epsilon}_{\theta}$ to the current $\boldsymbol{x}_{t}$, and adding a Gaussian noise perturbation to obtain a cleaner $\boldsymbol{x}_{t-1}$.
-Classifier-guided DDIM sampling [5] aims to restore images from noise conditioned on the class label.
-Given the diffusion model $\boldsymbol{\epsilon}_{\theta}$, the latent code $\boldsymbol{x}_{t}$ at timestep $t$, the classifier $p_{\theta}(y|x_{t})$, and the gradient scale $s$, the sampling process for obtaining $\boldsymbol{x}_{t-1}$ is formulated as:
-$\hat{\epsilon}=\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t})-\sqrt{\bar{\alpha}_{t-1}}\triangledown_{\boldsymbol{x}_{t}}\log p_{\phi}(y|\boldsymbol{x}_{t}),$ (2)
+The DDPM forward process, in which Gaussian noise is progressively added to the latent code of a clean image: $\bm{x}_{0}$:
+$\bm{x}_{t}=\sqrt{\alpha_{t}}\bm{x}_{0}+\sqrt{1-\alpha_{t}}\bm{z},$ (1)
+where $\bm{z}\sim\mathcal{N}(0,\mathbf{I})$ and $\left\{\alpha_{t}\right\}$ are the noise schedule.
+The backward process is aimed at gradually denoising $\bm{x_{T}}$, where at each step a cleaner image is obtained.
+This process is achieved by a U-Net $\bm{\epsilon}_{\theta}$ that predicts the added noise $\bm{z}$.
+Each step of the backward process consists of applying $\bm{\epsilon}_{\theta}$ to the current $\bm{x}_{t}$, and adding a Gaussian noise perturbation to obtain a cleaner $\bm{x}_{t-1}$.
+Classifier-guided DDIM sampling [5] aims to generate images from noise conditioned on the class label.
+Given the diffusion model $\bm{\epsilon}_{\theta}$, the latent code $\bm{x}_{t}$ at timestep $t$, the classifier $p_{\theta}(y|x_{t})$, and the gradient scale $s$, the sampling process for obtaining $\bm{x}_{t-1}$ is formulated as:
+$\hat{\epsilon}=\bm{\epsilon}_{\theta}(\bm{x}_{t})-\sqrt{\bar{\alpha}_{t-1}}\triangledown_{\bm{x}_{t}}\log p_{\phi}(y|\bm{x}_{t}),$ (2)
 and
-$\boldsymbol{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\cdot\frac{\boldsymbol{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{\epsilon}}{\sqrt{\bar{\alpha}_{t}}}+\sqrt{1-\bar{\alpha}_{t-1}}\hat{\epsilon},$ (3)
+$\bm{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\cdot\frac{\bm{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{\epsilon}}{\sqrt{\bar{\alpha}_{t}}}+\sqrt{1-\bar{\alpha}_{t-1}}\hat{\epsilon},$ (3)
 where $\alpha$ is the denoise schedule.
 In the self-attention block of the U-Net, features are projected into queries $\mathit{\mathbf{Q}}$, keys $\mathit{\mathbf{K}}$, and values $\mathit{\mathbf{V}}$.
-The output of the block $\boldsymbol{o}$ is obtained by:
-$\boldsymbol{o}=\mathbf{A}\mathbf{V},\;\;\textup{where}\;\mathbf{A}=\textup{Softmax}(\mathit{\mathbf{Q}}\mathit{\mathbf{K}}^{\top})$ (4)
-The self-attention operation allows for long-range interactions between features.
+The output of the block $\bm{o}$ is obtained by:
+$\bm{o}=\mathbf{A}\mathbf{V},\;\;\textup{where}\;\mathbf{A}=\textup{Softmax}(\mathit{\mathbf{Q}}\mathit{\mathbf{K}}^{\top})$ (4)
+The self-attention operation allows for long-range interactions between image tokens.
 3.2 Overview
-In generating infinite scenes, the process involves two key steps.
-Firstly, the RGB for the forthcoming frame is calculated based on the RGBD data and camera extrinsics of the current frame.
-Secondly, to address the presence of black holes in the next frame’s image, inpainting [34, 22, 48] and outpainting [16, 46] techniques are employed for correction.
-Despite this approach, several challenges arise: 1) The irregular nature of the black holes makes it difficult for existing inpainting methods to fill these areas while maintaining semantic coherence with the rest of the image.
-2) Inaccuracies in the depth data can cause distortions in certain image areas when wrapping.
-These distortions tend to accumulate over time with the movement of the camera, leading to a noticeable decline in image quality.
-3) Additionally, as the image advances, it requires the enhancement of details in certain areas, a task at which current inpainting and outpainting methods are inadequate, particularly due to their lack of super-resolution capabilities.
-To avoid generating distorted areas, ensure semantic consistency between adjacent frames, and enhance image detail during camera movement, we leverage the robustness and rich semantic capabilities of the denoise phase and decoder in the stable diffusion model.
-We edit the features of the diffusion based on depth and camera extrinsics, followed by denoising to produce a high-definition image for the subsequent frame.
-Our comprehensive design is detailed in Fig.
+Figure 2: Overview of our proposed pipeline.
+Starting from a real or generated RGBD ($I$, $D$) image at the current view, we apply DDIM inversion to obtain intermediate latent code $x_{t_{1}}$ at timestep $t_{1}$ using a pre-trained U-Net model.
+A warping with the high-pass filter strategy is applied to generate latent code for the next novel view.
+A few more DDPM forward steps from timestep $t_{1}$ to $t_{2}$ are applied for enlarging the degree of freedom w.r.t.
+the warped latent code.
+In the denoising process, we apply pre-trained U-Net to generate the novel view from $x_{t_{2}}^{\prime}$.
+The cross-view self-attention module and feature-correspondence guidance are applied to maintain the geometry correspondence between $x_{t_{2}}$ and $x_{t_{2}}^{\prime}$.
+The right side shows the warped image and our generated novel view $I^{\prime}$.
+Our method greatly alleviates blurring, inconsistency, and distortion.
+The overall pipeline is zero-shot and training-free.
+Perpetual view generation as the camera moves presents a complex challenge.
+This process involves seamlessly filling in unseen regions caused by image warping, adding details to objects as they come closer, while ensuring the imagery remains realistic and diverse.
+Prior works [18, 20, 3] have focused on training a refiner to enhance details and create new content for areas requiring inpainting or outpainting.
+These efforts have shown promising outcomes, yet the effectiveness of the refiner is generally limited to scenarios that align with the training dataset.
+Since diffusion models can generate high-quality large-variety images from random latent code, a direct solution arises: can we modify the powerful pre-trained text-to-image diffusion model as a refiner?
+Empirically, DDIM inversion strategy [14, 43] can obtain the intermediate latent code at each timestep and the image can be reconstructed by those latent codes.
+To this end, we attempt to explicitly warp the latent code of the current view and generate the novel view by the pre-trained text-to-image diffusion model.
+Our overall pipeline is illustrated in Fig.
 2.
-Initially, we obtain the latent code of the previous frame’s RGB image at timestep $t_{1}$ through the inverse diffusion process.
-We then wrap the current frame’s latent code using depth information and camera extrinsics.
-To maintain content and high-frequency detail consistency while wrapping the latent code, we introduce a low-pass wrapping strategy.
-Then several DDPM (Denoising Diffusion Probabilistic Models) processes are applied to increase the generative process’s flexibility, adding noise from timestep $t_{1}$ to $t_{2}$.
-Finally, we obtain the RGB image of the next frame through the denoising process.
-For ensuring semantic consistency between adjacent frames, we utilize a cross-frame self-attention strategy, injecting the key and value from the previous frame’s UNet network into the next frame.
-Furthermore, we propose a feature-correspondence-guidance strategy to influence and maintain semantic consistency between adjacent frames during the denoise latent code process.
-Our entire pipeline requires only a pre-trained text-to-image diffusion model and a depth estimation model, eliminating the need for any additional training or fine-tuning.
-3.3 Wrapping latent codes
-Algorithm 1 Low-pass latent code wrapping
-$\boldsymbol{x}_{t}$$\triangleright$ latent code at timestep $t$ of current view $c$
-1:$F(\boldsymbol{x}_{t})\leftarrow FFT(\boldsymbol{x}_{t})$triangleright Apply Fast Fourier Transform
-2:Split $F(\boldsymbol{x}_{t})$ into $F_{low}$ and $F_{high}$ using threshold $\sigma$
-3:$\boldsymbol{x}_{t}^{low}\leftarrow IFFT(F_{low})$triangleright Inverse FFT on low-frequency component
-4:$\boldsymbol{x}_{t}^{low-wrapped}\leftarrow Wrap(\boldsymbol{x}_{t}^{low})$triangleright Wrap the low-frequency content
-5:$F_{wrapped}\leftarrow FFT(\boldsymbol{x}_{t}^{low-wrapped})$triangleright FFT on wrapped content
-6:$F^{\prime}\leftarrow F_{wrapped}+F_{high}$triangleright Combine low-frequency of wrapped content with high-frequency of original code
-[…]
-[…]
+Initially, we obtain the latent code $x_{t_{1}}$ of the current view’s RGB image $I$ at timestep $t_{1}$ through the DDIM inversion process.
+We then warp the current frame’s latent code $x_{t_{1}}$ to the next view $x^{\prime}_{t_{1}}$ using depth information and camera extrinsic parameters.
+However, directly denoising from $x^{\prime}_{t_{1}}$ to the image also suffers from blurry, which results in the non-integer pixel coordinates and the interpolation operation.
+More noise is added from $x^{\prime}_{t_{1}}$ at timestep $t_{1}$ to $x^{\prime}_{t_{2}}$ at timestep $t_{2}$ by DDPM forward operation, for generating high-quality images.
+The side effect of DDPM is the geometry inconsistency between adjacent views.
+To this end, we propose a feature-correspondence-guidance denoising strategy to enforce geometry consistency.
+Moreover, a high-pass filtering strategy is proposed to maintain the consistency of high-frequency details between adjacent views.
+Please refer to Fig.
+3 for the motivation of our proposed modules.
+Our overall pipeline requires only a pre-trained text-to-image diffusion model and a depth estimation model, eliminating the need for any additional training or fine-tuning.
+3.3 Warping latent codes
+Algorithm 1 Warping latent code with high-pass filter
+1:$\bm{x}_{t}$triangleright latent code at timestep $t$ of current view $c$
+2:$F(\bm{x}_{t})\leftarrow FFT(\bm{x}_{t})$triangleright Apply Fast Fourier Transform
+3:Split $F(\bm{x}_{t})$ into $F_{low}$ and $F_{high}$ using threshold $\sigma$
+4:$\bm{x}_{t}^{low}\leftarrow IFFT(F_{low})$triangleright Inverse FFT on low-frequency component
+5:$\bm{x}_{t}^{low-warped}\leftarrow warp(\bm{x}_{t}^{low})$triangleright warp the low-frequency content
+6:$F_{warped}\leftarrow FFT(\bm{x}_{t}^{low-warped})$triangleright FFT on warped content
+7:$F^{\prime}\leftarrow F_{warped}+F_{high}$triangleright Combine low-frequency of warped content with high-frequency of original content
+8:$\bm{x}_{t}^{\prime}\leftarrow IFFT(F^{\prime})$triangleright Inverse FFT to get latent code for next view $c^{\prime}$ return $\bm{x}_{t}^{\prime}$triangleright warped latent code at timestep $t$ for next view $c^{\prime}$
 The results in the right side of Fig.
-2 reveal that directly wrapping images based on camera intrinsics $K$, extrinsics ($R,t$), and depth information leads to regions of distortion in the images.
-Additionally, the use of inpainting [34, 22, 48] and outpainting [16, 46] models to fill these gaps does not achieve satisfactory outcomes.
-In pursuit of generating photo-realistic images, and considering the robustness of the denoising process and decoder in the stable diffusion model, we opt to edit the latent code corresponding to timestep $t$.
+2 reveal that directly warping images based on camera intrinsics $K$, extrinsics $\left\{\mathcal{R},\mathcal{T}\right\}$, and depth information leads to regions of distortion in the images.
+Additionally, the use of inpainting [36, 23, 53] and outpainting [17, 49, 51] models to fill these gaps does not achieve satisfactory outcomes.
+In pursuit of photo-realistic images, we opt to edit the latent code corresponding to timestep $t$.
 PnP [44] and DIFT [42] have shown that the features of diffusion possess strong semantic information, with semantic parts being shared across images at each step.
-The simplest method for wrapping the latent code follows the same approach as wrapping the image.
-The only difference between wrapping the latent code and wrapping the image is a slight modification in the camera intrinsics; this entails scaling the camera intrinsics proportionally based on the different resolutions of the image and latent code.
-The overall procedure for wrapping the latent code is illustrated in Alg.
+The simplest method for warping the latent code follows the same approach as warping the image.
+The only difference between warping the latent code and warping the image is a slight modification in the camera intrinsics; this entails scaling the camera intrinsics proportionally based on the different resolutions of the image and latent code.
+The overall procedure for warping the latent code is illustrated in Alg.
 1.
-Initially, a latent code ${x}_{t}$ is obtained and transformed via Fast Fourier Transform (FFT) to $F(\boldsymbol{x}_{t})$.
-This is divided into low-frequency $F_{\text{low}}$ and high-frequency $F_{\text{high}}$ components, segregated at threshold $\lambda$.
-The key step involves wrapping the Inverse FFT (IFFT) processed low-frequency component $x_{t}^{\text{low}}=\text{IFFT}(F_{\text{low}})$, forming $\boldsymbol{x}_{t}^{low-wrapped}$.
-Merging $\text{FFT}(\boldsymbol{x}_{t}^{low-wrapped})$ with $F_{\text{high}}$, we obtain $F_{\text{new}}$, from which the final latent code $\boldsymbol{x}_{t}^{\prime}=\text{IFFT}(F^{\prime})$ is reconstructed.
+Initially, a latent code ${x}_{t}$ is obtained and transformed via Fast Fourier Transform (FFT) to $F(\bm{x}_{t})$.
+This is divided into low-frequency $F_{\text{low}}$ and high-frequency $F_{\text{high}}$ components, segregated at threshold $\sigma$.
+The key step involves warping the Inverse FFT (IFFT) processed low-frequency component $\bm{x}_{t}^{\text{low}}=\text{IFFT}(F_{\text{low}})$, warping to the next view $\bm{x}_{t}^{low-warped}$.
+Merging $\text{FFT}(\bm{x}_{t}^{low-warped})$ with $F_{\text{high}}$, we obtain $F^{\prime}$, from which the final latent code $\bm{x}_{t}^{\prime}=\text{IFFT}(F^{\prime})$ is reconstructed.
 This approach efficiently preserves high-frequency details, enabling high-fidelity scene generation aligned with text prompts.
@@
 Cross-view self-attention.
-To maintain consistency between the generated result and the original image, inspired by recent image and video editing works [39, 45, 8, 42], we modify the process of the self-attention module of U-Net when denoising the latent code $\boldsymbol{x}_{t}^{\prime}$.
+To maintain consistency between the generated result and the original image, inspired by recent image and video editing works [44, 48, 8, 45], we modify the process of the self-attention module of U-Net when denoising the latent code $\bm{x}_{t}^{\prime}$.
 Specifically, we denoise the views for the current and next view together.
@@
 The modified cross-view self-attention for generating a novel view is defined as:
-$\boldsymbol{o}^{\prime}=\mathbf{A^{\prime}}\mathbf{V},\;\;\textup{where}\;\mathbf{A}^{\prime}=\textup{Softmax}(\mathbf{Q}^{\prime}\mathbf{K}^{\top}),$ (5)
-where $\mathbf{Q}^{\prime}$, $\mathbf{A}^{\prime}$, and $\boldsymbol{o}^{\prime}$ are query, attention matrix, and output features for the novel views.
+$\bm{o}^{\prime}=\mathbf{A^{\prime}}\mathbf{V},\;\;\textup{where}\;\mathbf{A}^{\prime}=\textup{Softmax}(\mathbf{Q}^{\prime}\mathbf{K}^{\top}),$ (5)
+where $\mathbf{Q}^{\prime}$, $\mathbf{A}^{\prime}$, and $\bm{o}^{\prime}$ are query, attention matrix, and output features for the novel views.
 $\mathbf{K}$ and $\mathbf{V}$ are injected keys and values obtained from the self-attention module for generating the original view.
+Please note that the $\mathbf{K}$ and $\mathbf{V}$ are also warped before injection.
 Feature-correspondence guidance.
 Maintaining geometry consistency between adjacent views using the cross-view self-attention mechanism presents challenges, especially in preserving high-frequency details as the camera moves forward.
-The recent DIFT study [38] highlights the potential of using intermediate features of diffusion models for accurate point-to-point image matching.
+The recent DIFT [42] highlights the potential of using intermediate features of diffusion models for accurate point-to-point image matching [32].
 Additionally, the concept of vanilla classifier guidance [5] steers the diffusion sampling process using pre-trained classifier gradients towards specific class labels.
-Building on these ideas, we integrate feature correspondence guidance into the DDIM process to enhance consistency between adjacent views, addressing the challenge of detail preservation in dynamic scenes.
-Specifically, we obtain the features of the current and next view at each timestep $t$ of the DDIM process and calculate the cosine distance between the wrapped original features and features from the next novel views:
-$\mathcal{L}_{sim}^{t}=\frac{1-\cos\left[\mathrm{Wrap}(f_{t}),f_{t}^{\prime}\right]}{2},$ (6)
-where $f_{t}$ and $f_{t}^{\prime}$ are diffusion features extracted from pre-trained U-Net $\boldsymbol{\epsilon}_{\theta}$ at timestep $t$ and $\mathrm{Wrap}$ is the wrapping functions.
+Building on these ideas, we integrate feature correspondence guidance into the DDIM sampling process to enhance consistency between adjacent views, addressing the challenge of detail preservation in dynamic scenes.
+Specifically, we obtain the features of the current and next view at each timestep $t$ of the DDIM process and calculate the cosine distance between the warped original features and features from the next novel views:
+$\mathcal{L}_{sim}^{t}=\frac{1-\cos\left[\mathrm{warp}(f_{t}),f_{t}^{\prime}\right]}{2},$ (6)
+where $f_{t}$ and $f_{t}^{\prime}$ are intermediate features extracted from pre-trained U-Net $\bm{\epsilon}_{\theta}$ at timestep $t$ and $\mathrm{warp}$ is the warping functions.
 The lower $\mathcal{L}_{sim}^{t}$, the higher the similarity.
-We further introduce the similarity score $\mathcal{L}_{sim}$ to the DDIM sampling process, for generating novel views with geometry consistency. The predicted noise $\hat{\epsilon}$ is formulated as:
-$\hat{\epsilon}=\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t})-\lambda\sqrt{\bar{\alpha}_{t-1}}\triangledown_{\boldsymbol{x}_{t}}\mathcal{L}_{sim}^{t},$ (7)
-where $\lambda$ is the constant hyper-parameter and latent code $\boldsymbol{x}_{t-1}$ is calculated by Eq. 3
+We further introduce the similarity score $\mathcal{L}_{sim}^{t}$ to the DDIM sampling process, for generating novel views with geometry consistency.
+The predicted noise $\hat{\epsilon}$ is formulated as:
+$\hat{\epsilon}=\bm{\epsilon}_{\theta}(\bm{x}_{t})-\lambda\sqrt{\bar{\alpha}_{t-1}}\triangledown_{\bm{x}_{t}}\mathcal{L}_{sim}^{t},$ (7)
+where $\lambda$ is the constant hyper-parameter and latent code $\bm{x}_{t-1}$ is calculated by Eq. 3
 4 Experiments
-Figure 3: Qualitative comparisons of InfNat-0 [17] and ours.
-We provide four scene images with various styles and categories as start points and ask models to fly through the images.
-10 frames are illustrated for each starting point.
-Implementation details.
-We take Stable Diffusion [34] with the pre-trained weights from version 2.1https://huggingface.co/stabilityai/stable-diffusion-2-1-base as the basic text-to-image diffusion and MiDas [32] with weights $\mathrm{dpt\_beit\_large\_512}$https://github.com/isl-org/MiDaS.
-The overall diffusion timesteps is $1000$.
-We wrap the latent code at timestep $t_{1}$=$21$ and add more degrees of noise to timestep $t_{2}$=$441$.
-The threshold $\sigma$ for low-pass wrapping is $20$ and the hyper-parameter $\lambda$ for feature-correspondence guidance is $300$.
-Due to the page limit, please refer to the supplementary material (supp.) for details.
-Baselines.
-We compare against 1) 2 supervised methods for perpetual view generation: InfNat [19] and InfNat-0 [17].
-2) one text-conditioned scene generation: SceneScape [7].
-3) 2 supervised methods for text-to-video generation: CogVideo [11] and VideoFusion [23].
+4.1 Implementation details
+[…]
+4.2 Baselines
+We compare against 1) two supervised methods for perpetual view generation: InfNat [20] and InfNat-0 [18].
+2) one text-conditioned 3D point cloud-based scene generation: SceneScape [7].
+3) two supervised methods for text-to-video generation: CogVideo [11] and VideoFusion [24].
 4) one method for zero-shot text-to-video generation: T2V-0 [14].
-Evaluation Metrics.
-The perpetual view generation is evaluated both quantitatively and qualitatively.
-We evaluate CLIP [31] score, which indicates text-scene alignment for quantitative comparisons.
+4.3 Evaluation metrics
+We evaluate our zero-shot perpetual scene generation into two aspects: 1) the quality of generated images and text-image alignment, and 2) the temporal consistency of generated image sequences.
+Image quality and text-image alignment.
+We evaluate CLIP score [33], which indicates text-scene alignment for quantitative comparisons.
+A high average CLIP score indicates not only that the generated images are more aligned with the corresponding prompts but also that they consistently maintain high quality [14].
 CogVideo [11], VideoFusion [24], SceneScape [7], and T2V-0 [14] are all engaged in text-conditioned generation tasks.
-We generated 50 scene-related text prompts using GPT-4https://openai.com/gpt-4 and then created videos using each of the three methods.
+We generated 50 scene-related text prompts using GPT-4333https://openai.com/gpt-4 and then created videos using each of the three methods.
 For the InfNat [20] and InfNat-0 [18] methods, we used Stable Diffusion to generate the initial frame, followed by subsequent frame generation based on this initial frame.
 We calculated the distance between each generated frame and the text embedding, known as the CLIP score.
-A high average CLIP score indicates not only that the generated images are more aligned with the corresponding prompts but also that they consistently maintain high quality.
-Considering that the InfNat [19] and InfNat-0 [17] methods might gradually generate other scenes, we provided 10 very general prompts such as ‘an image of the landscape’ and ‘an image of the mountain’ for these methods, and then selected the highest CLIP score as the CLIP score for the current frame.
-4.1 Quantitative comparison
-Figure 4: Qualitative comparisons of SceneScape [7], T2V-0 [14], and our DreamDrone.
-We visualize 20 continuous frames for each textual prompt.
-As the camera flies, our method generates geometry-consistent scene sequences.
-The quantitative comparison results are shown in Tab.
+Considering that the InfNat [20] and InfNat-0 [18] methods trained on natural scene datasets, we further provided 10 very general prompts such as ‘an image of the landscape’ and ‘an image of the mountain’ for these methods, and then selected the highest CLIP score as the CLIP score for the current frame.
+Table 1: Ablations of image quality and temporal coherence of generated image sequences with various lengths. Please refer to Fig. 3 for quality comparisons.
+Methods PSNR $\uparrow$ SSIM $\uparrow$ CLIP $\uparrow$
+8 frames 16 frames 32 frames 8 frames 16 frames 32 frames 8 frames 16 frames 32 frames
+warp image 26.90 22.46 21.62 0.25 0.23 0.24 0.138 0.112 0.106
+warp latent 28.35 28.57 28.75 0.27 0.28 0.24 0.135 0.122 0.125
+warp latent+DDPM 24.67 23.04 22.59 0.12 0.10 0.06 0.302 0.297 0.308
+warp latent+DDPM+guidance 28.27 28.21 28.10 0.34 0.30 0.26 0.317 0.316 0.313
+warp latent+DDPM+guidance +cross-view attn. 28.89 28.83 28.75 0.32 0.31 0.27 0.318 0.315 0.315
+warp latent+DDPM+guidance +cross-view attn.+high pass filter 29.91 29.86 29.79 0.39 0.38 0.35 0.320 0.318 0.319
+Temporal consistency of generated image sequences.
+We demonstrate our advancements in temporal consistency against other SOTA methods by calculating average PSNR and SSIM scores across adjacent frames for generated videos with different lengths.
+The higher scores demonstrate the superiority in terms of cross-view consistency.
+4.4 Ablation studies
+We perform ablation studies on our three proposed modules: 1) warping latent with high-pass filter, 2) cross-view self-attention module, and 3) feature-correspondence guidance.
+The quantitative ablation results are shown in Tab.
+1 and we visualize the ablation samples in Fig.
+3.
+Figure 3: Ablation results for the key components.
+We perform ablation studies by disabling the key components of our method.
+We illustrate every five frames for each ablation experiment.
+Please zoom in for better comparisons.
+The simplest method for infinite scene generation tasks is frame-by-frame image warping, but this approach is unfeasible, as is directly warping the latent code.
+Warping images leads to non-integer pixel coordinates, resulting in interpolation-induced blurring and distortion.
+Moreover, these errors accumulate with each frame generated, leading to a collapse in quality.
+The first two rows of Table 1 show that directly warping images or warping latent codes (i.e., removing DDPM) results in very low CLIP scores, indicating poor quality of the generated images.
+The generated images becoming progressively blurred can also be observed in the first two rows of Fig.
+3.
+We introduce DDPM to increase the degrees of freedom of the diffusion model, thereby generating high-quality images.
+However, the introduction of DDPM has the side effect of worsening the semantic consistency between adjacent frames ($3^{rd}$ row in Tab.
+1 and Fig.
+3).With the help of DDPM, the CLIP score increases from $0.125$ to $0.308$ when generating 32 images.
+Please refer to the supp.
+for the generated results with different scales of the DDPM forward process.
+To ensure the quality of image generation while also maintaining consistency with adjacent views, we propose a feature-correspondence guidance strategy.
+Comparing the third and fourth rows of Fig.
+3, it is evident that the semantic consistency between adjacent frames is significantly enhanced after adding guidance, with noticeable improvements in both PSNR and SSIM scores in Tab.
 1.
-It is evident that the CLIP scores for the InfNat and InfNat-0 methods are low, indicating poor robustness.
-These methods are not sufficiently versatile to handle infinite scene generation for any scenario, as also illustrated by the visual results in Fig.
-3.
-CogVideo and VideoFusion, which are supervised text-to-video generation methods, struggle to generate first-person perspective landscapes due to limitations of text-video training datasets.
-Additionally, their CLIP scores decrease with an increasing number of frames, such as 32 frames, which is corroborated by the visual results in supp.
-T2V-0 [14], SceneScape [7], and our method, both based on the pre-trained stable diffusion model and training-free, have relatively higher CLIP scores.
-However, as the number of frames increases, SceneScape and T2V-0’s CLIP scores decrease, whereas ours remains constant (please refer to Fig.
-4 for qualitative comparisons).
-We attribute this to our low-pass wrapping strategy, which provides better content guidance for the generated images, and the introduction of feature-correspondence guidance, ensuring the coherence of adjacent frames and thus maintaining a higher CLIP score.
-Table 1: Quantitative comparison results. We calculate the average CLIP score [31] across various video lengths.
-8 frames 16 frames 32 frames
-InfNat [19] 0.125 0.123 0.118
-InfNat-0 [17] 0.128 0.125 0.122
-SceneScape [7] 0.263 0.238 0.174
-CogVideo [11] 0.255 0.249 0.241
-VideoFusion [23] 0.281 0.283 0.272
-T2V-0 [14] 0.312 0.305 0.287
-Ours 0.320 0.318 0.319
+To further enhance cross-view consistency, we adopted cross-view attention modules and high-pass filtering.
+From the visualized results at the $5^{th}$ and $6^{th}$ rows in Fig.
+3, it is clear that the semantic consistency of adjacent camera perspectives is further strengthened after incorporating the cross-view attention module.
+The operation of the high-pass filter further preserves the high-frequency details of the current frame, thereby further enhancing the semantic consistency of high-frequency details between adjacent frames.
+For instance, comparing the left side house at the $4^{th}$, $5^{th}$, and $6^{th}$ rows in Fig.
+3, the cross-view consistency is enhanced after adding the proposed modules.
+In addition to conducting ablation experiments on the modules we propose, we continue to explore two more questions:
+Figure 4: Ablation study for scene travel.
+We visualize two image sequences and change the prompt when generating novel views.
+We illustrate every five images and the prompts are changed when generating $31^{th}$ image ($7^{th}$ image shown in each row).
+Q1: Can DreamDrone shuttle from one scene to another by changing text prompts?
+During the frame-by-frame process, we changed the textual prompts, with the generated results shown in Fig.
+4.
+The visualized results demonstrate that DreamDrone can smoothly complete the scene travel (from streets in Inazuma City to urban art street) or the transition of scene styles (from realistic to Lego style) while ensuring the semantic consistency of adjacent views, according to the changes in textual prompts.
+Figure 5: Ablation study on customized camera trajectory.
+We generate images with different camera directions.
+For the sample of the Eiffel Tower, our camera perspective continuously ascends.
+For the $2^{nd}$ scene of the Lego city, our camera not only moves forward but also shifts upwards and to the right.
+Q2: Can explicitly warping the latent code control the trajectory of camera perspective movement?
+Since our method generates image sequences frame by frame, we can freely adjust the camera’s flight angle by altering the camera’s extrinsic parameters.
+In Fig.
+5, we provide sequences of images generated under different camera trajectories.
+The results show that our method possesses a high degree of freedom, allowing for the free customization of the camera’s trajectory.
+Other state-of-the-art methods cannot achieve this functionality.
 4.5 Qualitative comparison
+Figure 6: Qualitative comparisons of InfNat-0 [18] and ours.
+We provide four starting scene images with various styles and categories as start points and ask models to fly through the images.
+50 frames are generated and we illustrate every five frames for each starting scene image.
 In our comparison with InfNat-0 [18] (Fig.
-3), focusing on various scenes including coastlines, rivers, Van Gogh-style landscapes, and city streetscapes, we identified four main differences: Firstly, InfNat-0 shows proficiency in coastline scenes, a reflection of its training data, but our training-free DreamDrone surpasses it in later frames due to InfNat-0’s cumulative errors over time.
+6), focusing on various scenes including coastlines, rivers, Van Gogh-style landscapes, and city streetscapes, we identified four main differences: Firstly, InfNat-0 shows proficiency in coastline scenes, a reflection of its training data, but our training-free DreamDrone surpasses it in later frames due to InfNat-0’s cumulative errors over time.
 Secondly, in natural scenes with closer objects, InfNat-0’s flawed generation becomes more apparent, whereas our method maintains consistency.
@@
 However, as seen in Fig.
-4, both methods have limitations.
+7, both methods have limitations.
 SceneScape struggles with outdoor scenes and forward camera movement, leading to blurred and distorted results after 8 steps due to its reliance on a pre-trained inpainting model.
-T2V-0 displays a decline in quality beyond the third frame in complex environments like Lego-style cities, likely from its latent code editing approach that compromises frame continuity and geometric consistency.
+T2V-0 displays a drop in quality beyond the third frame in complex environments like Lego-style cities, likely from its latent code editing approach that compromises frame continuity and geometric consistency.
 Conversely, our DreamDrone excels across various scenes.
@@
 Please refer to supp.
-for more quantitative comparisons.
-Our task bears similarities to text-to-video generation, we further provide qualitative comparisons with VideoFusion [23], one of the state-of-the-art methods for text-to-video generation.
-Due to the page limit, please refer to supp.
-for detailed comparisons.
-4.3 Ablation studies
-Figure 5: Ablation study on the key components of our method.
-We perform ablation studies by disabling the key components of our method: feature-correspondence guidance, low-pass wrapping strategy, and cross-view self-attention module.
-We perform ablation studies on 1) low-pass feature wrapping, 2) cross-view self-attention module, and 3) feature-correspondence guidance in Fig.
-5.
-Without feature-correspondence guidance and cross-view attention, even if we perform the corresponding wrapping at the feature level, it is challenging to ensure strong consistency between generated adjacent frames ($1^{st}$ row).
-Direct wrapping of the latent code leads to increasingly blurred images and a loss of high-frequency details.
-However, the use of our low-pass wrapping strategy effectively ensures that the generated images retain high-definition details ($2^{nd}$ row).
-Without the guidance of the feature-correspondence guidance mechanism or the presence of a cross-view self-attention module, it is difficult to guarantee the geometric consistency of adjacent frames ($3^{rd}$ and $4^{th}$ rows).
-The components in our proposed pipeline enhance each other, ensuring that each generated frame is not only rich in detail but also maintains good consistency with adjacent frames.
-Please refer to supp.
-for more detailed ablation studies.
+for more comparisons.
+Figure 7: Qualitative comparisons of SceneScape [7], T2V-0 [14], and our DreamDrone.
+We visualize 20 continuous frames for each textual prompt.
+As the camera flies, our method generates geometry-consistent scene sequences.
+As our task bears similarities to text-to-video generation, we further provide qualitative comparisons with VideoFusion [24]. Due to the page limit, please refer to supp. for detailed comparisons.
+4.6 Quantitative comparison
+Tab.
+2 offers a detailed comparison of various SOTA methods for generating image sequences, including our method, DreamDrone.
+When compared to other training-based methods, DreamDrone, despite being training-free, consistently achieves higher CLIP scores across all frame lengths (0.320, 0.318, 0.319 for 8, 16, and 32 frames respectively).
+This is particularly noteworthy as the CLIP scores for training-based methods generally degrade as the number of generated frames increases.
+For instance, VideoFusion’s [24] CLIP scores decrease from 0.281 for 8 frames to 0.272 for 32 frames.
+This trend suggests a decline in the quality of generated images with an increase in sequence length for training-based methods.
+Table 2: Qualitative comparisons with other SOTA methods. We evaluate the quality and temporal coherence of the generated image sequences with various lengths.
+Methods PSNR $\uparrow$ SSIM $\uparrow$ CLIP $\uparrow$
+8 frames 16 frames 32 frames 8 frames 16 frames 32 frames 8 frames 16 frames 32 frames
+training-based InfNat [20] 28.75 28.67 28.65 0.32 0.30 0.30 0.125 0.123 0.118
+InfNat-0 [18] 28.92 28.89 28.87 0.37 0.35 0.34 0.128 0.125 0.122
+CogVideo [11] 31.03 30.08 29.32 0.45 0.39 0.31 0.255 0.249 0.241
+VideoFusion [24] 29.89 28.36 28.78 0.41 0.37 0.31 0.281 0.283 0.272
+training-free T2V-0 [14] 27.25 26.17 26.03 0.27 0.24 0.23 0.312 0.305 0.287
+Scenescape [7] 29.87 29.75 29.66 0.41 0.38 0.34 0.318 0.282 0.279
+DreamDrone (Ours) 29.91 29.86 29.79 0.39 0.38 0.35 0.320 0.318 0.319
+In contrast, DreamDrone maintains high CLIP scores even as the sequence length increases, indicating superior image quality.
+When compared to other training-free methods, DreamDrone also stands out.
+For example, while T2V-0’s [14] CLIP scores decrease from 0.312 for 8 frames to 0.287 for 32 frames, DreamDrone’s CLIP scores remain relatively stable, further demonstrating its robustness in maintaining image quality across varying sequence lengths.
+This analysis underscores the effectiveness of DreamDrone in generating high-quality, temporally coherent image sequences without the need for training.
 5 Conclusion
-In this work, we pioneer the concept of zero-shot, training-free text-to-perpetual view generation, introducing DreamDrone for generating geometry-consistent scenes.
-We harness the robust capabilities of a pre-trained text-to-image diffusion model and develop a novel feature-correspondence-guidance diffusion process for creating novel views.
-Our quantitative and qualitative analyses underscore DreamDrone’s superiority over other state-of-the-art methods.
-We anticipate that our approach will inspire further innovation in AI-Generated Content (AIGC) and make advanced research accessible to a broader spectrum of researchers.
-Supplementary Material
+In this work, we propose DreamDrone, a novel approach for generating flythrough scenes from textual prompts without the need for training or fine-tuning.
+Our method explicitly warps the intermediate latent code of a pre-trained text-to-image diffusion model, enhancing the quality of the generated images and the generalization ability.
+We propose a feature-correspondence-guidance diffusion process and a high-pass filtering strategy to ensure geometric and high-frequency detail consistency.
+Experimental results indicate that DreamDrone surpasses current methods in terms of visual quality and authenticity of the generated scenes.
+Acknowledgement
+This project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Award Number: MOE-T2EP20122-0006).
+DreamDrone: Text-to-Image Diffusion Models
+are Zero-shot Perpetual View Generators
+Supplement Material
+[…]
+In this supplementary material, we provide more comprehensive ablation studies, comparisons of visual results with text-to-video methods, and additional visual results of our method.
+Finally, we discuss the limitations and social impact of this approach.
 6 Implementation details
-We take Stable Diffusion [34] with the pre-trained weights from version 2.1https://huggingface.co/stabilityai/stable-diffusion-2-1-base as the basic text-to-image diffusion and MiDas [32] with weights $\mathrm{dpt\_beit\_large\_512}$https://github.com/isl-org/MiDaS.
-The overall diffusion timesteps is $1000$.
-We wrap the latent code at timestep $t_{1}$=$21$ and add more degrees of noise to timestep $t_{2}$=$441$.
-The threshold $\sigma$ for low-pass wrapping is $20$ and the hyper-parameter $\lambda$ for feature-correspondence guidance is $300$.
-We conducted the experiments on Titan-RTX GPU.
-The generated speed is roughly 30 seconds per frame.
+[…]
 7 Ablation studies
-In this section, we conduct our ablation studies on each proposed module.
-First, we explore the impact of wrapping the latent code at timestep $t_{1}$ in our pipeline on the generation results, with comparative results shown in Fig.
-6.
-According to the principles of diffusion, a smaller timestep means that the corresponding latent code has less noise, thus retaining more complete content and detailed information.
-When we wrap the latent code at a smaller timestep (for example,$t_{1}=1$ or $21$), it is observed that, with the camera’s movement, DreamDrone still maintains its complete content information quite well.
-As the timestep increases, with more noise, the content information gets somewhat disrupted, and wrapping the latent code then does not yield as favorable results.
-This experiment demonstrates that the latent code corresponding to a smaller timestep possesses richer geometry and content information.
-Figure 6: Ablation study on the wrapping timestep $t_{1}$.
-Next, we investigate how editing the latent code in different ways can enhance the generation results.
-We set various thresholds for the low-pass wrapping module, where a smaller threshold indicates the retention of less low-frequency information and more high-frequency information.
-The comparative results are detailed in Fig.
-7.
-The first and second rows display the generation results corresponding to smaller thresholds.
-It can be observed that setting an excessively small threshold retains too much ineffective or noisy high-frequency information, leading to generated images with an unrealistic abundance of high-frequency details.
-Conversely, setting the threshold too high, or removing the low-pass wrapping module, results in a lack of retained detail and very blurry images.
-Figure 7: Ablation study on the threshold of the low-pass wrapping module $\sigma$.
-After obtaining the wrapped latent code, we next investigate the impact of setting $t_{2}$ on the generation results.
-The size of $t_{2}$ influences the freedom in the denoising process; a larger $t_{2}$ means adding more noise to the wrapped latent code, thereby gaining higher degrees of freedom.
-The generation results corresponding to different $t_{2}$ settings are shown in Fig.
+Figure 8: Ablation results for the key components.
+We perform ablation studies by disabling the key components of our method.
+We illustrate every five frames for each ablation experiment.
+Please zoom in for better comparisons.
+In this section, we first provide additional ablation results for Fig.
+8 as referenced from Fig.
+3 in the main text.
+To enhance the robustness of our ablations, we include one example for each experimental setup.
+The comprehensive results of the ablation study for each component are illustrated in Fig.
 8.
-A smaller $t_{2}$ (for example, $t_{2}=161$ or 241) results in insufficient noise added to the wrapped latent code, leading to increasingly blurred generated images.
-Setting a larger $t_{2}$ produces clearer images with richer details, but an excessively high degree of freedom can decrease the geometric consistency between adjacent frames, adversely affecting their coherence.
-Figure 8: Ablation study on the denoising timestep $t_{2}$.
-To ensure geometric consistency between adjacent frames during the denoising process, we introduced a feature-correspondence guidance strategy.
-Different guidance weights lead to significant variations in generation results, as detailed in Fig.
-9.
-When feature-correspondence guidance is not introduced (i.e., $\lambda=0$), the generated image sequence maintains content consistency, but some details may not be consistent.
-The introduction of feature-correspondence guidance ensures the consistency of details between adjacent frames; however, excessively high guidance weights can impact the realism of the generated images.
-Figure 9: Ablation study on the weight parameter of feature-correspondence guidance $\lambda$.
+For quantitative results, please refer to Tab.
+1 in the main text.
+The simplest method for tasks involving infinite scene generation is frame-by-frame image warping.
+However, this approach is impractical, as is the direct warping of the latent code.
+Warping images results in non-integer pixel coordinates, which leads to interpolation-induced blurring and distortion.
+Furthermore, these errors accumulate with each generated frame, causing a significant degradation in quality, with the images becoming progressively blurred, as shown in warp image and warp latent in Fig.
+8.
+To address these challenges, we introduce DDPM to increase the degrees of freedom for the diffusion model, facilitating the generation of high-quality images (warp latent + DDPM).
+With the incorporation of DDPM, the CLIP score improves from $0.125$ to $0.308$ for a series of 32 images.
+Nevertheless, the introduction of DDPM inadvertently affects the semantic consistency between adjacent frames.
+To maintain the integrity of image content, and inspired by previous methods [44, 14], we integrate cross-view attention modules into our framework.
+As demonstrated in warp latent + DDPM + cross attn.
+in Fig.
+8, the consistency of geometries across views is significantly improved compared to warp latent + DDPM.
+To ensure high-quality image generation while also maintaining consistency across adjacent views, we propose a feature-correspondence guidance strategy.
+Comparing warp latent + DDPM + guidance with warp latent + DDPM + guidance + cross view attn., it is evident that semantic consistency between adjacent frames is significantly enhanced after incorporating guidance, as indicated by the noticeable improvements in both PSNR and SSIM scores in Table 1 of the main text.
+To further improve the cross-view consistency of high-frequency details, we have employed high-pass filtering.
+This approach aids in preserving the high-frequency details of the current frame, thus enhancing the semantic consistency of high-frequency details between consecutive frames.
+For instance, the enhanced cross-view consistency, particularly of the house on the left side in Fig.
+8, illustrates the effectiveness of adding the proposed modules.
+Then, we conduct more detailed ablation studies on each proposed module in a Q&A manner.
+Q1:
+Does DDIM inversion limit the reconstruction fidelity?
+Figure 9: DDIM inversion and image reconstruction.
+We illustrate the pipeline for evaluating the reconstruction performance using DDIM inversion.
+The left side is the pipeline and the right side is the reconstructed results at different timesteps $t_{1}$.
+We visualize two different result samples.
+Before warping latent, the most important thing is to ensure we can reconstruct the original image without any editing of the intermediate latent code.
+To this end, we establish a simple experiment.
+We obtain the intermediate latent code $x_{t_{1}}$ at different timestep $t_{1}$, denoising the noise, and decode the reconstructed image.
+In  Fig.
+9, the left side is the pipeline of this experiment and the right side is the reconstructed images at different timestep $t_{1}$.
+We take two sample images as examples.
+From the results, we can figure out that the images can be reconstructed at different timestep $t_{1}$.
+Based on the former discussion, regarding the top branch in Fig.
+2 in the main text, the image at the current view can be reconstructed.
+The reconstruction of the top branch is the foundation of the feature-correspondence guidance and cross-view attention.
+Q2:
+Why do the image sequences become blurrier when generating more images, no matter whether warping the image or warping the latent code?
+The $1^{st}$ and $2^{nd}$ rows in Fig.
+3 in the main text show that the images become blurrier when generating more images.
+Besides, the reconstruction results in  Fig.
+9 show that DDIM inversion can reconstruct original images if there is no editing for the intermediate latent code, i.e., warping.
+It is straightforward that when we fly through, in other words, zoom in, the images, the images will become much blurrier.
+This is because the warping operation leads to non-integer pixel coordinates.
+Previous SOTA methods [20, 18, 3] train a refiner to add the details and inpainting or outpainting the missing region when the camera is moving.
+In this paper, we serve the pre-trained text-to-image diffusion model as a ‘refiner’ due to its powerful generation capacity.
+Q3:
+Why DDPM forward is needed?
+Figure 10: Ablation studies for DDPM forward without high-pass filtering.
+To evaluate the necessity of DDPM forward operation, we conduct ablation experiments based on the simple pipeline.
+The corresponding ablation results are shown in Fig.
+11
+Now we analyze the necessity of the DDPM forward process.
+As illustrated in Fig.
+2 in the main text, we further apply DDPM forward after warping the latent code.
+Comparing the $5^{th}$ and $6^{th}$ rows in Fig.
+8, we can figure out that the image quality improves a lot after DDPM is applied.
+The details are enhanced and there is no distortion.
+The side effect of DDPM forward is that the correlation between adjacent views degrades because more degrees of freedom are introduced by DDPM.
+Figure 11: The visualization results for evaluating DDPM forward without high-pass filtering.
+We illustrate the results based on the pipeline shown in Fig.
+10.
+We fix the warping timestep $t_{1}$ and generate image sequences with different DDPM forward timesteps $t_{2}$.
+To facilitate the comparisons, we further illustrate the generation results using our overall pipeline at last.
+We visualize every five frames per sample.
+To evaluate how DDPM affects the generation results, we fix the warping timestep $t_{1}$ and illustrate the generated image sequences with DDPM at various timesteps $t_{2}$.
+The pipeline of this ablation experiment is shown in Fig.
+10.
+The results are shown in Fig.
+11.
+A smaller $t_{2}$ means less degree of freedom for the diffusion model, which can result in blurring and distortion.
+As shown in the first two rows in Fig.
+11, the generated images become blurrier when generating more images.
+A proper $t_{2}$ makes the geometry between adjacent views more consistent.
+In the $3^{rd}$ and $4^{rd}$ rows in Fig.
+11, the geometry layout in the image sequences becomes consistent.
+For instance, the geometry of the house on the left side of the image looks roughly consistent.
+Moreover, the image quality is satisfied.
+As $t_{2}$ becomes larger, more random noise are added to the warped latent code.
+Though the image quality is promising, the consistency degrades a lot.
+As shown in the $5^{th}$ and $6^{th}$ rows, the consistency across adjacent views is much worse than $3^{th}$ and $4^{th}$ rows.
+Figure 12: Ablation studies for high-pass filtering without DDPM forward.
+To analyze if the DDPM can be replaced by the high-pass filter, we conduct a simple ablation based on this pipeline.
+The corresponding ablation results are shown in Fig.
+13.
+This ablation demonstrates that the DDPM forward module with a proper timestep $t_{2}$ improves the image quality.
+But the consistency is still not satisfied.
+That’s the reason why we further propose the feature-correspondence guidance strategy.
+Moreover, since the high-pass filter preserves details from the previous view, is it possible to remove DDPM forward and only use the high-pass filter?
+To this end, we further conduct an ablation experiment.
+The pipeline for this ablation is shown in Fig.
+12.
+In this pipeline, we remove DDPM forward operation and add the high-pass filter when warping the latent code.
+The experimental results are shown in Fig.
+13.
+We show two results for each $\sigma$.
+No matter how large the threshold $\sigma$ is, the high-pass filter cannot help to preserve the details from previous view.
+The reason is that the high-pass filter can only preserve high-frequency details from the previous view, rather than the low-frequency content.
+Combined with the results shown in Fig.
+14, the low-frequency information dominants the content when combining frequencies from different images.
+As discussed before, the content would become much blurrier when warping the latent code, which motivates us to propose the feature-correspondence guidance strategy.
+Figure 13: The visualization results for high-pass filter without DDPM forward.
+To evaluate if we can preserve high quality from the previous view using the high-pass filter, we remove the DDPM forward module and visualize the image sequences with different threshold $\sigma$.
+To facilitate the comparisons, we further illustrate the generation results using our overall pipeline at last.
+We visualize every five frames per sample.
+Q4:
+Will the combination of low and high frequencies from different images break up the correlation of the frequency of the original image and introduce more errors?
+Figure 14: Toy experiments of low and high-frequency combination from different images.
+Our toy experiments are illustrated on the left side.
+We combine the low frequency of Elon Musk and the high frequency of Vincent van Gogh’s self-portrait with various threshold $\sigma$.
+The higher $\sigma$, the more low-frequency of Elon Musk is used.
+The results with various $\sigma$ are illustrated on the right side.
+Please zoom in for comparisons.
+To evaluate the feasibility of the frequency combination, we conduct a toy experiment, which is shown in Fig.
+14.
+In this experiment, we first obtain the frequency from two different images and combine the frequencies given different threshold $\sigma$.
+As shown on the right side in Fig.
+14, the content of Elon Musk does not change much with different $\sigma$, which demonstrates the feasibility of frequency combination.
+An extremely small $\sigma$, for instance, $\sigma=10$, introduces excessive details from van Gogh’s portrait.
+In this toy experiment, the content of the two images is extremely different.
+However, regarding the perpetual view generation task, the content of the adjacent view would not be so different.
+Now we analyze how different $\sigma$ affects the generation results.
+We apply all the proposed modules in this experiment and change the $\sigma$ value.
+The comparison results are shown in Fig.
+15.
+As shown in Fig.
+15, a small $\sigma$ neglects more low-frequency content from the previous view, which results in an inconsistency between images.
+A large $\sigma$ introduces less high-frequency details from the previous view.
+Though looks consistent, the generated images look not as realistic as $\sigma=20$.
+Figure 15: Ablation studies on high-pass filter. We apply all the proposed modules with various thresholds $\sigma$ of the high-pass filter.
 8 Additional qualitative comparisons
-Figure 10: Qualitative comparisons of VideoFusion [23] and our DreamDrone.
+Figure 16: Qualitative comparisons of VideoFusion [24] and our DreamDrone. We show the visualization results given three prompts and illustrate every five frames for each sample.
 Our task bears similarities to text-to-video generation, with the key difference being that text-to-video generation cannot be controlled by camera pose, and the quality significantly diminishes as the number of generated frames increases.
 VideoFusion [24], one of the state-of-the-art methods for video generation tasks, has been visually compared with our method, which is illustrated in Fig.
-10.
+16.
 It is evident that VideoFusion’s generated results become blurry with an increase in frame count, and the effect of camera movement is less pronounced.
@@
 9 More visualization results
-In this section, we provide more visualization results.
-Please refer to Figs.
-11 and 12 for details.
-Moreover, please watch our attached videos for a comprehensive evaluation of our proposed DreamDrone.
+In this section, we provide more visualization results. We generate 120 images for each prompt and visualize one image from every third frame. Please refer to Figs. 17 and 18 for details.
 Figure 17: Visualization results of our DreamDrone.
+We generated 120 image sequences for each text prompt and visualized one image from every third frame to demonstrate the model’s capability in producing diverse and stable visual outputs over time.
 Figure 18: Visualization results of our DreamDrone.
+We generated 120 image sequences for each text prompt and visualized one image from every third frame to demonstrate the model’s capability in producing diverse and stable visual outputs over time.
 10 Limitation
@@
 We plan to address these shortcomings in our future work.
+11 Social impact
+We introduce a new method for creating perpetual scenes from text descriptions, making it easier for people to generate high-quality images without needing complex training or data.
+This breakthrough can help in various areas, such as making educational content more engaging, aiding in environmental planning, and giving creative professionals new tools to express their ideas.
+As this technology becomes available, it’s important to use it wisely, ensuring it benefits society and does not contribute to misinformation or unethical use.
+In essence, DreamDrone offers exciting possibilities for innovation while emphasizing the need for responsible use.